Find out Ka for 10^-2M HCN having pOH=10.............................................

Grade:12th pass

1 Answers

Khimraj
6 years ago
The dissociation of KCN can be written as
HCN(aq) + H2O(l) \rightleftharpoons H3O+(aq) +CN-(aq)
Ka = [H3O+][CN-]/[HCN]
pOH = 10 So pH = 14 – 10 = 4
pH = -log[H3O+]
-4 = log[H3O+]
[H3O+] = 10-4
[CN-] =[H3O+] = 10-4
So Ka = (10-4*10-4)/(10-2) = 10-6.
Hope it clears if you like answer then please approve it.